OpenStudy (anonymous):

John is standing in front of a tree at point S. He can see an arc of the tree trunk that stretches from J to K measuring 64°. What is the measure of angle JSK? 58° 116° 148° 32°

OpenStudy (anonymous):

May someone please help me?

OpenStudy (anonymous):

Pretty please?

OpenStudy (ganpat):

<JSK = 1/2 (arc Jk ) = (64)/2 = 32 degrees.. if my formula is right.. i guess it is right .. :-p

OpenStudy (anonymous):

THANK YOU!!!!

OpenStudy (anonymous):

@DarlingDare LOL I have seen you ask many questions, and every time you say, "May someone please help me?" Well... IT'S DRIVING ME CRAZY! It is grammatically incorrect!!!